OpenStudy (anonymous):

Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

switch around the equations.... \[2x=3y-11\] \[2x=9-y\] substitute one into the other, since both equal 2x \[3y-11=9-y\] simplify \[4y=20\] solve for y \[y=5\] plug in y into one equation \[3(5)-2x=11\] solve for x \[-2x=-4\] answer: \[x=2\] \[y=5\]